OpenStudy (anonymous):

A point M on a segment with endpoints X (1,-2) and Y (10,3) partition the Segment in 4:1 ratio. find M.

OpenStudy (anonymous):

omg marry me!! :DDD

OpenStudy (haseeb96):

M(x,y)= (m1x2-m2x1)/m1 +m2 , (m1y2 - m2y1)/m1+m2 let (x1,y1)=(1,-2) (x2,y2) =(10,3) m1:m2=4:1 now just put it in the formula and get the answer

OpenStudy (anonymous):

could you tell me the answer and explain how u did it

OpenStudy (bradely):

x= (m1x2+m2x1)/(m1+m2) y= (m1y2+m2y1)/(m1+m2)

OpenStudy (haseeb96):

M(x,y)= (m1x2+m2x1)/m1 +m2 , (m1y2 + m2y1)/m1+m2 sorry this is the correct formula M(x,y)= ( (4)(10) + (1)(1)/ 4+1 , ( (4)(3) +(1)(-2) ) / 4+1 = 40 +1 /5 , 12 -2 /5 = 41/5 ,10/5 M(x,y) = (41/5 ,2) this is the point of M

OpenStudy (haseeb96):

@Samidancer16 if u have any difficulty then ask i have solved it first i wrote the wrong formula but now i have corrected it

OpenStudy (bradely):

x= (4*10+1*1)/(4+1)=41/5 y= (4*3+1*-2)/(4+1)=10/5

OpenStudy (anonymous):

thank uou that helped so much!!

OpenStudy (haseeb96):

Not say thanks just help three persons anywhere and say to them that help three more persons anywhere